A nurse has 1 liter of a solution that is 20% alcohol. How much pure alchol must she add to bring the solution up to a 25% concentration
Answer by stanbon(75887)

A nurse has 1 liter of a solution that is 20% alcohol. How much pure alcohol must she add to bring the solution up to a 25% concentration
----------------
20% solution DATA;
Amt. = 1 liter ; amt of active ingredient = 0.20x liter
----------------
100% alcohol DATA;
Amt. = x liter; amt of active ingredient = x liter
----------------
25% Mixture DATA;
Amt. = (1+x) liter ; amt of active ingredient = 0.25(1+x) = 0.25+0.25x liter
----------------------
EQUATION:
active + active = active
0.20x + x = 0.25+0.25x
1.2x = 0.25 + 0.25x
0.95x = 0.25
x = 0.263 liter (amount of pure alcohol that must be added)
==================
Cheers,